What is the solubility of BaSO4 in 0.10M of Na2SO4 solution? How do i solve this
Answer #1

BaSO4 is not that solube so

BaSO4 <-> Ba+2 + SO4-2

Ksp = [Ba+2][SO4-2]

from databooks

Ksp = 1.1*10^-10

and we know

[SO4-2] = 0.1 (from the Na2SO4 solution)

so

Ksp = [Ba+2][SO4-2]

substitute

[Ba+2] = Ksp/[SO4.-2] = (1.1*10^-10)/0.1 = 1.1*10^-9 M

since S = Ba+2 then

S = 1.1*10^-9 M